OpenStudy (precalchelp):

using the derivative rule find the rate of change 8x^-7 - 6x^-2 - x^-5

OpenStudy (aum):

\(f(x) = 8x^{-7} - 6x^{-2} - x^{-5}\) Rate of change of f(x) is f'(x). Just differentiate f(x) with respect to x.

OpenStudy (aum):

\[\large \frac{d}{dx}x^n = nx^{n-1} \\ \text{For example} \\ \frac{d}{dx}x^{-7} = (-7)x^{-7-1} = -7x^{-8} \\ \]